A 0.07 mol calcium reacts with 0.04 mol oxygen gas to produce calcium oxide, CaO. Given that calcium is the first to be completely consumed, calculate the expected mass of CaO produced in this experiment


1
Expert's answer
2021-10-26T05:01:01-0400

2Ca + O2 → 2CaO


Moles of CaO = 0.04 × 2 = 0.08


Mass of CaO = 56.0774 × 0.08=4.486grams